OpenStudy (olivya999):

Find the inverse of the function. f(x) = 7x + 2

OpenStudy (sparrow2):

if you want i cant teach how to deal with is and you will do it by yourself @olivya999

OpenStudy (olivya999):

f-1(x) = quantity x plus two divided by seven Not a one-to-one function f-1(x) = quantity x minus two divided by seven f-1(x) = x divided by seven minus two

OpenStudy (olivya999):

okay

OpenStudy (sparrow2):

f(x)=7x+2 and y=7x+2 is the same f(x)=y

OpenStudy (sparrow2):

then you switch places y goes to x and x goes to y : you will have x=7y+2

RhondaSommer (rhondasommer):

f(x) = 7x + 2 replace f(x)(basically y) with x x=7y+2 so solve to get y alone

OpenStudy (sparrow2):

then you try to have y seperated, here it is y=(x-2)/7

OpenStudy (olivya999):

so c?

OpenStudy (sparrow2):

yep

OpenStudy (olivya999):

thanks @sparrow2

RhondaSommer (rhondasommer):

@sparrow2 you should know by now that any direct answer is not allowed

OpenStudy (sparrow2):

and there did i write direct answer? @RhondaSommer

OpenStudy (sparrow2):

*where

RhondaSommer (rhondasommer):

"then you try to have y seperated, \[here-it-is-y=(x-2)/7\]"

OpenStudy (sparrow2):

i explaind the whole process and she understood .i see no problem here

RhondaSommer (rhondasommer):

she never said that. she just asked if the answer was c. AFTER you gave the answer to her